Need help[ filling the boxes!

Lott Company uses a job order cost system and applies overhead to production on the basis of direct labor costs. On January 1, 2017, Job No. 50 was the only job in process. The costs incurred prior to January 1 on this job were as follows: direct materials $24,600, direct labor $14,760, and manufacturing overhead $19,680. As of January 1, Job No. 49 had been completed at a cost of $110,700 and was part of finished goods inventory. There was a $18,450 balance in the Raw Materials Inventory account.

During the month of January, Lott Company began production on Jobs 51 and 52, and completed Jobs 50 and 51. Jobs 49 and 50 were also sold on account during the month for $150,060 and $194,340, respectively. The following additional events occurred during the month.

1. Purchased additional raw materials of 110,700 on account.

2. Incurred factory labor cost of $86,000. Of this amount $19,680 related to employer payroll taxes.

3. Incurred manufacturing overhead costs as follows: indirect materials $20,910; indirect labor $24,600; depreciation expense on equipment $14,760; and carious other manufacturing overhead costs on account $19,680.

4. Assigned direct materials and direct labor to jobs as follows.

Job No. 50 51 52 Direct Materials Direct Labor $12,300 47,970 36,900 $6,150 30,750 24,600

Open job cost sheets for Jobs 50, 51, and 52. Enter the January 1 balances on the job cost sheet for Job No. 50. Job No. 50 wain Beg Direct Materials Direct Labor Manufacturing Overhead Jan Cost of completed job Direct materials Direct labor Manufacturing overhead Total cost

Job No. 51 Date Jan Direct Materials Direct Labor Manufacturing Overhead Direct materials Iirercs kibor PM:aruuiacuriocticatd Total cost

Job No. 52 Date Direct Materials Direct Labor Manufacturing Overhead Jan
Answer #1

Answer:

Lott Company uses a job order cost system and applies overhead to production on the basis of direct labor costs.

Calculation of Overhead rates:

Manufacturing overhead cost = Indirect materials + indirect labor + depreciation expense on equipment + other manufacturing overhead costs

= $20,910 + $24,600 + $14,760 + $19,680

= $79,950

To direct labor cost = $6,150 + 30,750 + 24,600 = $61,500

Applied Overhead rate = 79,950 / 61500 = $1.30 per direct labor dollar

Hence:

Applied Manufacturing Overhead for Jan:

Job 50 = 6150 * 1.30 = $7,995

Job 51 = 30750 * 1.30 = $30,975

Job 52 = 24600* 1.30 = $31,980

Job No. 50 Date Beg Jan. Direct Material Direct Labor Manufacturing Overhead $ 19,680 $7,995 $27,675 $14,760 24,600 12,300 $ 6,150 36,900 20,910 Cost of completed Job Direct materials Direct Labor Manufacturing overhead Total cost $ 36,900 $20,910 $ 27,675 85,485

Job No. 51 Date Beg Jan. Direct Material Direct Labor Manufacturing Overhead 47,970 $30,750 47,970 30,750 $39,975 $ 39,975 Co

Job No. 52 Date Jan Direct Materia Direct Labor Manufacturing Overhead $31,980 kr 36,900 $ 24,600